OpenStudy (anonymous):

can someone PLEASE please please pleeease help me! I'll reward you with a fan and medal. which of the following has the greatest x value for the vertex? explain your choice. A. f(x) = x^2 +4x + 4 B. f(x) = 3(x-1)^2+5 C. f(x) = -2(x+2)^2+3 D. f(x) = -3x^2+6x+1

OpenStudy (jango_in_dtown):

A. f(x)=x^2+4x+4=(x+2)^2 vertex is (-2,0) B.f(x)=3(x-1)^2+5 vertex is (1,5) C.f(x)=-2(x+2)^2+3 vertex is (-2,3) D.f(x) = -3x^2+6x+1=-3(x^2-2x-1)=-3(x-1)^2+6 vertex is (1,6)

OpenStudy (jango_in_dtown):

@rosamartinez

OpenStudy (anonymous):

so, A? @jango_IN_DTOWN

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

the x-coordinate of the vertex of a generic parabola: \(y=ax^2+bx+c\) is: \[x = - \frac{b}{{2a}}\]

OpenStudy (anonymous):

I have no idea how to do this, honestly.

OpenStudy (michele_laino):

for example, if we consider the parabola in A) we get: \[x = - \frac{b}{{2a}} = - \frac{4}{{2 \cdot 1}} = - 2\] being: \(a=1, b=4\)

OpenStudy (michele_laino):

now, I consider the parabola in B). I can rewrite such parabola as below: \[\begin{gathered} y = 3{\left( {x - 1} \right)^2} + 5 = 3\left( {{x^2} + 1 - 2x} \right) + 5 = \hfill \\ \hfill \\ = 3{x^2} + 3 - 6x + 5 = 3{x^2} - 6x + 8 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

I have computed the square of the binomial: \((x-1)^2\) so we get: \[y = 3{x^2} - 6x + 8\] as we can see, by comparison with the generic formula of a parabola, we have: \(a=3,b=-6,c=8\) so, using my formula above, we get: \[x = - \frac{b}{{2a}} = - \frac{{ - 6}}{{2 \cdot 3}} = 1\] namely the vertex of that parabola is located at \(x=1\)

OpenStudy (michele_laino):

please do the same with the parabola D.

OpenStudy (michele_laino):

hint: there we have: \(a=-3,b=6,c=1\) so: \[x = - \frac{b}{{2a}} = - \frac{6}{{2 \cdot \left( { - 3} \right)}} = ...?\]

OpenStudy (michele_laino):

please complete

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

it is simple: here is the answer: \[x = - \frac{b}{{2a}} = - \frac{6}{{2 \cdot \left( { - 3} \right)}} = 1\]

OpenStudy (anonymous):

so is it A, B, C or D? give an example of a perfect square trinomial and the difference of squares binomial, the show how to factor. 144r^2 + 24r + 1 5 5^2 25 A difference of squares can be factored using the following pattern: a2−b2=(a+b)(a−b) @Michele_Laino

OpenStudy (michele_laino):

an example of the perfect square trinomial, can be this: \(x^2+49+14x=(x+7)^2\)

OpenStudy (michele_laino):

the pattern is: \((a+b)^2=a^2+b^2+2 a b\)

OpenStudy (michele_laino):

for example, if we have to compute the square of this binomial: \((12r+1)^2\) we have \(a=12r, b=1\) so, if we apply the pattern above, we can write: \(a^2+b^2+2 a b = (12r)^2+ (1)^2+ 2 \cdot 12 r \cdot 1=...\) please complete

OpenStudy (michele_laino):

hint: \((12r)^2=12r \cdot 12r= (12 \cdot 12) r^2=...?\)

OpenStudy (michele_laino):

by the way, for first question, the answer is B. and D.

OpenStudy (michele_laino):

what is \( 12 \cdot 12=...?\)

OpenStudy (anonymous):

144

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

so we have: \[{\left( {12r + 1} \right)^2} = 144{r^2} + 1 + 24r\] furthermore, a difference of square binomial, can be this: \(25x^2-49\)=\((5x-7)\cdot(5x+7)\)

OpenStudy (anonymous):

okay, so the first one is a square trinomial and the second a square binomial right? and for the first question it can only be one choice and I need to explain my choice. @Michele_Laino

OpenStudy (michele_laino):

more precisely, the first one: \((12r+1)^2=144r^2+1+24r\) is the square trinomial whereas the second one: \(25x^2-49=(5x-7)(5x+7)\) is the difference of squares

OpenStudy (michele_laino):

the vertex for each parabola is: \((-2,0)\) for parabola in A \((1,5)\) for parabola in B \((-2,3)\) for parabola in C \((1,4)\) for parabola in D

OpenStudy (michele_laino):

if we are looking for the vertex with the highest y-coordinate, then our answer is option B